OpenStudy (highschoolmom2010):

A friend says that there are 6720 different ways to combine 5 out of 8 ingredients to make a stew. Explain the error and find the correct answer.

ganeshie8 (ganeshie8):

you're "choosing" 5 ingredients out of available 8

ganeshie8 (ganeshie8):

so clearly "order" doesnt matter

ganeshie8 (ganeshie8):

that means u should use a "combinaiton", but ur friend is incorrectly using "permutation"

OpenStudy (highschoolmom2010):

i know that they used a permutation to get their answer but....i have tried using the combination but the answer that i am getting just isnt right :/

ganeshie8 (ganeshie8):

8c5 is the answer

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=8+choose+5

ganeshie8 (ganeshie8):

u can choose 5 things from 8 things in 8c5 ways

OpenStudy (highschoolmom2010):

i have to give the right answer too i think i got it \[\frac{ 8*7*6*5*4*3*2*1 }{ 3*2*1*5*4*3*2*1}=\frac{ 40320 }{ 720 }=56\]

OpenStudy (highschoolmom2010):

ah ha i got it :)

ganeshie8 (ganeshie8):

56 is \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

good job !

OpenStudy (highschoolmom2010):

thanks :)

ganeshie8 (ganeshie8):

np :) remember below for permutations :- u use permutation when "order of picking things matter". for example :- total possible names using letters "a, b, c, d, e"

ganeshie8 (ganeshie8):

remember below for combinations :- u use combination when "order of picking things DOESNT matter". for example :- present problem.

ganeshie8 (ganeshie8):

if u remember these two, u wud never get confuse wid which one to use when...